OpenStudy (anonymous):

Express the complex number in trigonometric form. -6i

OpenStudy (anonymous):

I know the form but I keep getting either 80 degrees or 260 degrees when I do arctan(6)

OpenStudy (anonymous):

z= a+bi in this case a=0 and b=-6 to find theta (tangtheta=b/a) and r=sq(a^2+b^2)

OpenStudy (anonymous):